OpenStudy (anonymous):

2 g of iron (III) oxide is reacted with 0.2 g of carbon according to: Fe2O3 + 3 C -> 2 Fe + 3 CO Determine the mass of iron (in g) produced.

OpenStudy (joannablackwelder):

Can you determine which one is limiting?

OpenStudy (anonymous):

So far I have calculated n(Fe2O3)=2/159.7=0.125 mol n(C)=0.2/12.01=0.0167 mol N(C)needed=0.0125mol(Fe2O3)x3mol(C)/1mol(Fe2O3)=0.0375 as there is only 0.0167molC available and 0.0375mol is needed then C is the limiting reagent. after that I used the molC available to determine the mass Fe produced: 0.-167x1/3x159.687/1=0.8889g but the answer was wrong so I dont think I am doing it right.

OpenStudy (joannablackwelder):

Carbon is limiting :)

OpenStudy (joannablackwelder):

Only the last step is the issue. The question is asking for grams of Fe, not iron oxide.

OpenStudy (anonymous):

Can you help me to determine where I am going wrong?

OpenStudy (joannablackwelder):

0.0167 mol C (2 mol Fe/3 mol C) * (55.8 g Fe/1 mol Fe)

OpenStudy (anonymous):

oh Duhh!!! Omg how did I not see that!

OpenStudy (joannablackwelder):

Haha, no worries. It happens to the best of us.

OpenStudy (joannablackwelder):

It would have been right if you were looking for iron oxide ;)

OpenStudy (joannablackwelder):

Welcome to Open Study!

OpenStudy (anonymous):

Yeah fail! Thank you so much!

OpenStudy (joannablackwelder):

No worries. Happy studies!

OpenStudy (joannablackwelder):

Oh, and if you are satisfied that a question has been answered, please award the answerer with a medal and close the question. Thanks!